OpenStudy (anonymous):

What is the slope of the line that contains the point (-2,8) (-2,1)

OpenStudy (anonymous):

Slope = (y2-y1) / (x2-x1)

OpenStudy (anonymous):

m is representing the slope here so for finding the slope m=y2-y1/x2-x1 since (x1,y1)=(-2,8) (x2,y2)=(-2,1) put these points in the above formula and get the slope and show it to me

OpenStudy (anonymous):

-3 , 10?

OpenStudy (anonymous):

what? m=1-8/-2+2 m=-7/0 m=infinite this could not exist